how do you sove this equation step by step (11-y)+ 0.5y=7.7 I need to solve for y.
Answer by mananth(16946)

(11-y)+ 0.5y=7.7
11-y +0.5y =7.7
-0.5y = 7.7-11
-0.5y = -3.3
y= -3.3/-0.5
y= 6.6